# ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images

### v3

@@ 3.2 Hand Pose Annotation
Figure 3: Automatic annotation pipeline of 3D hand pose. We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands [99] and SAM [100]. Various error terms are formulated to optimize the MANO parameters.
3.3 TherFormer: A Baseline Method
@@
Figure 4: Overall Framework of TherFormer. Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction. Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
@@
Mask-guided Spatial Transformer. Human hands are highly articulated objects that can adopt a wide variety of poses, often against complex backgrounds. We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction. Given backbone features, we first utilize a mask head to estimate the binary hand mask in the thermal image. Then, we leverage the deformable self-attention [101] to refine the hand spatial features under the guidance of the estimated hand mask. Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations. In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter. Lastly, we reduce the spatial dimensions of the spatial features with a series of convolutions layers for efficiency.
Temporal Transformer. Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities. To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention [102] to explicitly attend to the feature vector of every frame. The output is frame-wise spatio-temporal feature embeddings. Note that the temporal self-attention degrades to an MLP for single-image based setting.
In the pose head, we use the MLP to project the embeddings to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$. We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cfSec. 3.2). For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately. Please see more details of TherFormer in our supplementary document.
4 Experiment
4.1 Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec. 3.2) and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames. To that end, we first annotate the 2D joint locations from both egocentric and exocentric images and obtain the 3D joint positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
As shown in Tab. 2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of [43, 41, 37]. The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture. We also observe that only combining $\mathcal{L}_{mask}$ and $\mathcal{L}_{j2d}$ can already provide a plausible accuracy since they fit the projection of the 3D hand pose to two heterogeneous views. $\mathcal{L}_{mesh}$, though it hardly improves the joint accuracy, can result in more natural hand mesh. Adding $\mathcal{L}_{j3d}$ further refines the joints as it induces the explicit constraint to their positions. We showcase some annotation examples in Fig. 5. As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subjects’ hand color and shape.
Figure 5: Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
@@ 4.1 Evaluation of the Annotation Method
Table 2: Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
4.2 Experiment Setup
Dataset Preparation. We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments. We annotate the main part of our dataset (cfTab. 1) automatically following Sec. 3.2, of which the training and validation sets serve as the foundation for the training of all network models.
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT [15] and A2J-Transformer [46], and reproduced them on our dataset for experiments. HTT [15] is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer [46] only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines. We exclude the additional action block used by HTT [15], focusing solely on hand pose estimation. We adjusted the anchor initialization phase of A2J-Transformer [46] to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
Evaluation Metrics. We evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) [34], in both camera space and root-aligned (RA) space [38]. For RA space, we align the estimated wrist with its groundtruth position before measurement. For PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
4.3 Thermal Image-based 3D Hand Pose Estimation
Main Results. To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods [15, 46] on the main testing set, as shown in Tab. 3. TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT [15] given the same sequential images as input. Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec. 3.3). A performance gap can be observed between single image-based and video-based settings for both HTT [15] and TherFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices. Moreover, TherFormer-V also improves the performance over HTT [15] for other spectra (cfTab. 4), proving its adaptability to different inputs.
@@ 4.3 Thermal Image-based 3D Hand Pose Estimation
(a) A2J-Transformer [46] Single 51.68 0.474 20.76 0.603 34
(b) HTT [15] Single 49.09 0.489 20.69 0.599 211
@@
Table 3: Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
@@
Ablation Studies. We ablate our proposed mask-guided attention mechanism and the entire spatial transformer (cfSec. 3.3) to observe their impact. As can be seen in Tab. 3, the mask-guided attention mechanism notably contributes to TherFormer’s performance growth (row (c) vs. (e)), demonstrating its effectiveness in mitigating the effect of background clutter. Other components in the spatial transformer bring a large performance gain in the RA space (row (e) vs. (d)), confirming the importance of spatial feature enhancement for fine-grained hand joints localization.
4.4 Comparison between Spectrum
4.4.1 Qualitative Results under Challenging Conditions
@@
To justify the advantages of using thermal cameras for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on three of our auxiliary testing sets, i.e., gloves, darkness and sun glare (cf. Tab. 1). Our automatic annotation pipeline (cf. Sec. 3.2) becomes infeasible since hands appear corrupted in either RGB or depth images. Therefore, we conduct a qualitative analysis of their performance and present some representative examples in Fig. 6. We use the same model, i.e., HTT [15], trained on different spectra for a fair comparison. Please see the supplementary materials for more figure examples and demo videos under challenging conditions.
@@
RGB vs. Thermal. As can be seen, RGB-based methods fail with gloves wearing (cfFig. 6 (a-b)) and in darkness (cfFig. 6 (c-d)). Gloves change how human hands look and hide their natural colors and textures. Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process. Contrary to RGB techniques, as shown in Fig. 6 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
@@
Figure 6: Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
4.4.2 Quantitative Results under Normal Conditions

### v4

@@ 3.1. Multi-Spectral Hand Pose Dataset
Figure 4. Automatic annotation pipeline of 3D hand pose. We utilize the multi-view RGB and depth images as the input source and retrieve constraint information with off-the-shelf MediaPipe Hands (mediapipe_hands_2020, ) and SAM (kirillov2023segment, ). Various error terms are formulated to optimize the MANO parameters.
4. Hand Pose Annotation
To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in (hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ), to obtain the 3D hand pose ground truth for our dataset. In particular, we use the MANO statistical hand model (MANO, ) to represent 3D hand pose. The MANO model parameterizes the hand mesh vertices $\mathcal{V}(\beta,\theta)$ into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$, consisting of 45 parameters accounting for 15 hand joint angles (3 DoF for each) plus the rest for global rotation and translation. Following the MANO PyTorch version in (hasson19_obman, ), we denote the $N_{\mathcal{J}}=21$ hand joints mapped from the hand parameters as $\mathcal{J}(\beta,\theta)$. The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
(1) $\theta^{*}$ $=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$
$+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$
where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 4.
Initialization. For each sequence, we optimize the shape parameter $\beta$ [^5] For simplicity, we omit $\beta$ in Eq. 1 and subsequent equations. together with $\theta$ only at the first frame using Eq. 1 until convergence and fix its values for the subsequent frames. We initialize the pose parameter $\theta$ for each frame using the optimization result from the last frame. This helps to accelerate the convergence as well as keep the temporal consistency.
2D Joint Error $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}$. Given RGB images from $N_{C}$ viewpoints, we infer 2D hand joints $\mathcal{J}^{2D}$ with MediaPipe Hands (mediapipe_hands_2020, ) and define the 2D joint error as:
(2) $\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{2D}_{c,i}-\pi_{c}(\mathcal{J}(\theta)_{i})||$
where $\pi_{c}(\cdot)$ returns the 2D projection location for 3D position in the $c$-th camera viewpoint, and $\alpha_{c}$ is hyperparameter used to weigh different viewpoints.
Figure 5. Overall Framework of TherFormer. Backbone features are input to the mask-guided spatial transformer and temporal transformer to enhance the spatial representation and temporal interaction. Spatio-temporal embeddings are fed into the pose head to regress the 3D hand pose.
2D Mask Error $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}$. To generate the high-quality 2D hand mask, we prompt the prevalent Segment Anything Model (kirillov2023segment, ) with the 2D hand joints $\mathcal{J}^{2D}$ and the bounding box derived from it. We penalize the distance between the hand mesh vertices $\mathcal{V}(\theta)$ and the binary hand mask $\mathcal{M}_{c}$ as:
(3) $\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}=\sum_{c=1}^{N_{C}}\alpha_{c}\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{M}_{c,j}-\pi_{c}(\mathcal{V}(\theta)_{i})||$
where $\mathcal{M}_{c,j}$ is the coordinate of $j$-th non-zero pixel in the mask $\mathcal{M}_{c}$.
3D Joint Error $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}$. We triangulate the 2D joints from multiple views to lift them to 3D joints $\mathcal{J}^{3D}$ and measure their difference to $\mathcal{J}(\theta)$, which is written as:
(4) $\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}=\sum_{i=1}^{N_{\mathcal{J}}}||\mathcal{J}^{3D}_{i}-\mathcal{J}(\theta)_{i}||$
3D Mesh Error $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}$. To better supervise the hand mesh and fasten the optimization, we generate the 3D hand pose cloud $\mathcal{P}$ by cropping the depth image using the 2D hand mask and merging all views together. The 3D mesh error term compensates for the distance between the hand mesh and point cloud:
(5) $\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}=\sum_{i=1}^{N_{\mathcal{V}}}\min_{j}||\mathcal{P}_{j}-\mathcal{V}(\theta)_{i}||$
Regularization $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$. To alleviate irregular hand articulation, we constrain the joint angles to pre-defined lower and upper boundaries $\underline{\theta}$ and $\overline{\theta}$:
(6) $\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}=\sum_{i=1}^{45}(\max(\underline{\theta_{i}}-\theta_{i},0)+\max(\theta_{i}-\overline{\theta_{i}},0))$
Note that we also impose regularization to the shape parameter $\beta$ during initialization. For more implementation details of our dataset annotation process, please refer to Appendix B.
5. TherFormer: A Baseline Method
@@
Mask-guided Spatial Transformer. Human hands are highly articulated objects that can adopt a wide variety of poses, often against complex backgrounds. We propose a mask-guided spatial transformer module to accurately identify and focus on the intricacies of hand poses during spatial feature interaction. Given backbone features, we first utilize a mask head to estimate the binary hand mask in the thermal image. Then, we leverage the deformable self-attention (zhu2020deformable, ) to refine the hand spatial features under the guidance of the estimated hand mask. Specifically, we only take feature elements whose spatial locations are within the hand area as queries and sample keys from only the hand area and its surrounding locations. In this way, we not only reduce the computation waste on the irrelevant region but also increase the robustness to background clutter. Lastly, we reduce the spatial dimensions of the spatial features with a series of convolutions layers for efficiency.
Temporal Transformer. Temporal information is crucial for 3D hand pose estimation when coping with occlusion and solving ambiguities. To model temporal relationships, we first flatten the spatial features into 1D feature vectors and then employed the temporal self-attention (vaswani2017attention, ) to explicitly attend to the feature vector of every frame. The output is frame-wise spatio-temporal feature embeddings. Note that the temporal self-attention degrades to an MLP for single-image based setting.
In the pose head, we use the MLP to project the embeddings to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$. We leverage the binary cross-entropy loss to supervise the hand segmentation with the mask ground truth rendered from the annotated hand mesh (cfSec. 4). For 3D hand joint positions, we measure the L1 distance of its 2D projection and depth to that of the ground truth separately.
@@
Table 2. Evaluation of annotation results. The average 3D joint errors across all frames are reported (in cm). $\mathcal{L}_{reg}$ is used for all to mitigate irregular hand poses.
Figure 6. Examples of 3D hand pose annotations. Top row: left (blue) and right (red) hand 3D joints projected onto egocentric RGB images. Bottom row: visualization of hand mesh annotation.
@@
     (a)      A2J-Transformer (jiang2023a2j, )      Single      51.68      0.474      20.76      0.603      34
     (b)      HTT (wen2023hierarchical, )      Single      49.09      0.489      20.69      0.599      211
@@
Table 3. Comparison between TherFormer and state of the arts on thermal camera-based 3D hand pose estimation. The fps indicates the number of inference steps that models can run per second.
6. Experiment
6.1. Evaluation of the Annotation Method
As the first step, we validate the accuracy of our 3D hand pose annotation (cfSec. 4) and analyze the impact for optimization results. For evaluation, we manually annotate two random sequences from our main dataset, with a total of over 600 frames. To that end, we first annotate the 2D joint locations on RGB images from two viewpoints and obtain the 3D joint positions by triangulation. We calculate the average 3D joint errors across all frames to measure the accuracy.
As shown in Tab. 2, our annotation method achieves an average joint error of nearly 1cm, comparable to the results of (kwon2021h2o, ; hampali2020honnotate, ; zimmermann2019freihand, ). The multi-view setting shows remarkably better precision than the ego-view only optimization, demonstrating the necessity of multi-camera capture. We also observe that only combining $\mathcal{L}_{mask}$ and $\mathcal{L}_{j2d}$ can already provide a plausible accuracy since they fit the projection of the 3D hand pose to two heterogeneous views. $\mathcal{L}_{mesh}$, though it hardly improves the joint accuracy, can result in more natural hand mesh. Adding $\mathcal{L}_{j3d}$ further refines the joints as it induces the explicit constraint to their positions. We showcase some annotation examples in Fig. 6. As can be seen, both hand joint and mesh can be accurately annotated across different actions despite the presence of occlusion and the variance in subjects’ hand color and shape. Please see our project page for more visualization of annotation.
Figure 7. Impact of the temporal sequence length $T$ to TherFormer. The plots show all four metrics MEPE (mm) $\downarrow$, AUC $\uparrow$, MEPE-RA (mm) $\downarrow$, and AUC-RA $\uparrow$ against 5 sequence length settings, i.e., $T\in${1, 2, 4, 8, 16}.
6.2. Experiment Setup
Dataset Preparation. We utilize our own dataset for experiments as it uniquely contains egocentric images from multiple spectra, essentially for our benchmark experiments. We annotate the main part of our dataset (cfTab. 1) automatically following Sec. 4, of which the training and validation sets serve as the foundation for the training of all network models. Our automatic annotation pipeline (cf. Sec. 4) becomes infeasible under challenging scenarios, i.e., gloves, darkness and sun glare, since hands appear corrupted in either RGB or depth images. To facilitate the quantitative evaluation under challenging settings, we manually annotate the ground truth for a few sequences collected in the glove and sun glare scenarios (cfTab. 1). Specifically, we first annotate the 2D keypoints from two viewpoints and then use triangulation to obtain their 3D positions.
Method and Implementation. To provide sufficient baselines for follow-up works, we selected two state-of-the-art methods in 3D hand pose estimation: HTT (wen2023hierarchical, ) and A2J-Transformer (jiang2023a2j, ), and reproduced them on our dataset for experiments. HTT (wen2023hierarchical, ) is a video-based method thus enabling the evaluation in both two problem settings while A2J-Transformer (jiang2023a2j, ) only works under the single image-based setting. For a feasible comparison, we use the same sequence length, i.e., $T=8$, for HTT and TherFormer-V baselines. We exclude the additional action block used by HTT (wen2023hierarchical, ), focusing solely on hand pose estimation. We adjusted the anchor initialization phase of A2J-Transformer (jiang2023a2j, ) to better accommodate our dataset, without altering the density of its anchors. All trained models are tested on a single NVIDIA RTX 4090 GPU to fairly compare their inference speed.
Evaluation Metrics. Following HTT (wen2023hierarchical, ), we evaluate the accuracy of 3D hand pose estimation with two metrics: Percentage of Correct Keypoints (PCK) and Mean End-Point Error (MEPE) (zimmermann2017learning, ), in both camera space and root-aligned (RA) space. For the RA space, we align the estimated wrist with its groundtruth position before measuring the joint errors. In terms of PCK, we report the corresponding Area Under the Curve (AUC) over the 0-50mm/80mm error thresholds for the camera/RA space.
6.3. Thermal Image-based Results
Main Results. To assess TherFormer’s performance in thermal image-based 3D hand pose estimation, we compare it against state-of-the-art methods (wen2023hierarchical, ; jiang2023a2j, ) on the main testing set, as shown in Tab. 3. TherFormer-S outforms two competing methods under the single image-based setting, while TherFormer-V surpasses the counterpart HTT (wen2023hierarchical, ) given the same sequential images as input. Such an improvement mainly stems from our mask-guided spatial attention design that can adaptively encode the spatial interaction among hand joints with the guidance of the hand mask (cfSec. 5). A performance gap can be observed between single image-based and video-based settings for both HTT (wen2023hierarchical, ) and TherFormer. We credit this to their usage of temporal information that helps to tackle the occlusion cases and solve ambiguities. Thanks to our lightweight network design, TherFormer is highly efficient to run with fps of 136 and 52 for single and sequence input respectively, ensuring its real-time application to resource-constrained devices. Moreover, TherFormer-V also improves the performance over HTT (wen2023hierarchical, ) for other spectra (cfTab. 5), proving its adaptability to different inputs.
Figure 8. Examples of results with various spectra under three challenging settings. For visualization, we show the projection of the estimated left (red) and right (blue) 3D hand joints on images.
TherFormer-V (glove) TherFormer-V (sun glare)
Spectrum MEPE-RA (mm) $\downarrow$ AUC $\uparrow$ MEPE-RA (mm) $\downarrow$ AUC $\uparrow$
RGB 51.94 0.141 38.24 0.252
Depth 45.96 0.206 42.27 0.254
NIR 39.83 0.282 90.84 0.093
Thermal 39.23 0.302 32.56 0.363
Table 4. Comparison between different spectra under two challenging conditions, i.e., glove and sun glare.
HTT (Sequence) TherFormer-V TherFormer-S Best
RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.30 0.542
Depth 41.62 0.559 39.70 0.581 39.84 0.579 39.70 0.581
NIR 41.57 0.562 40.79 0.575 40.98 0.575 40.79 0.575
Thermal 47.07 0.512 43.36 0.549 44.64 0.539 43.36 0.549
Table 5. Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
@@ 6.3. Thermal Image-based Results
Ablation Studies. We ablate our proposed mask-guided attention mechanism and the entire spatial transformer to observe their impact. As can be seen in Tab. 3, the mask-guided attention mechanism notably contributes to TherFormer’s performance growth (row (c) vs. (e)), demonstrating its effectiveness in mitigating the effect of background clutter. Other components in the spatial transformer bring a large performance gain in the RA space (row (e) vs. (d)), confirming the importance of spatial feature enhancement for fine-grained hand joints localization.
Impact of Sequence Length. Here we conduct experiments to analyze the impact of the sequence length $T$ on the performance of our baseline method. We present the variation in our performance with respect to sequence length in Fig. 7. All experiments are conducted using the same hyperparameters, except for the $T=16$ model, which uses a smaller batch size due to GPU memory constraints. It can be observed that performance generally improves with increased sequence length, particularly in root-aligned metrics. There is a slight performance decrease in the camera space when the sequence length $T$ is increased to 16 but the performance in the root-aligned space is improved. The general trend underscores the importance of temporal information in enhancing the accuracy of 3D hand pose estimation.
6.4. Comparison between Spectrum
6.4.1. Qualitative Results under Challenging Conditions
@@
To justify the advantages of using thermal images for egocentric hand pose estimation under challenging scenarios, we conduct a comparison between four spectra on sequences collected in challenging scenarios (cf. Tab. 1). To intuitively show the results, we first conduct a qualitative analysis of their performance and present some representative examples in Fig. 8. We use the same model, i.e., HTT (wen2023hierarchical, ), trained on different spectra for a fair comparison. Please refer to our project page for more figure examples and demo videos under challenging conditions.
@@
RGB vs. Thermal. As can be seen, RGB-based methods fail with gloves wearing (cfFig. 8 (a-b)) and in darkness (cfFig. 8 (c-d)). Gloves change how human hands look and hide their natural colors and textures. Since RGB algorithms rely on skin’s texture and color to identify hand parts and joints, gloves, particularly those with solid colors or textures unlike skin, can interfere with this identification process. Contrarily, as shown in Fig. 8 (a-b), thermal imaging methods excel in identifying hands by leveraging the principles of heat conduction through gloves, effectively bypassing the limitations imposed by color and texture variations.
@@
6.4.2. Quantitative Results under Challenging Conditions
To provide the numerical results, we evaluate our TherFormer-V models trained for different spectra on our manually annotated sequences (cfSec. 6.2) and calculated the quantitative results presented in Tab. 4. As can be seen, thermal imaging-based approaches demonstrate the best performance among different spectra in challenging settings, underscoring thermal imagery’s advantages in difficult lighting conditions and when hands are occluded. Notably, the conclusion drawn from Tab. 4 resonate with the visualization in Fig. 8. The RGB spectrum performs poorly in scenarios involving gloves, where the color and texture of human hands are significantly altered. Intense sunlight introduces severe artifacts in NIR images, markedly impairing performance under sun glare.
6.4.3. Quantitative Results under Normal Conditions